How many molecules of carbon dioxide are present in a sample of (4.1x10^2) moles?


1
Expert's answer
2021-07-18T01:27:33-0400

1 mole = 6.023 × 1023 molecules

= 4.1×102 × 6.023 × 1023

= 2.47 × 1026 molecules